A cargo was dropped down an helicopter, which hovered at an height of h=125m .When will the cargo reach the land? g=10ms-².Air resistance must be ignored.


1
Expert's answer
2020-08-04T16:00:09-0400

The time of motion of the cargo

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 125\:\rm m}{10\:\rm m/s^2}}=5\:\rm s"